On a canoe trip, Rita paddled upstream (against the current) at an average speed of 2mi/h relative to the river bank. On the return trip downstream (with the current), her average speed was 3mi/h. Find Rita's paddling speed in still waters and the speed of the river's currents.
Answer by mananth(16946)

boat speed x Numbers
current speed ------------ y Numbers
x + 1 y = 3 .............1
x + -1 y = 2 .............2
Eliminate y
1 x + y = 3
1 x + - y = 2
Add the two equations
2 x = 5
/ 2
x = 2 1/2 = 2.50 mph boat speed
plug value of x in (1)
1 x + 1 y = 3
2.5 + 1 y = 3
1 y = 3 -2.5
1 y = 0.5
y = 1/2 mph